.Find all points (x, y) where the functions f(x), g(x), h(x) have the same value:
f(x) = 2^x−5 + 3, g(x) = 2x − 5, h(x) = 8/x+ 10
1
Expert's answer
2020-11-12T18:24:35-0500

Find all points (x, y) where the functions "f(x), g(x), h(x)" have the same value:


"f(x)=-2(x+5)+3, g(x)=2x-5, h(x)=\dfrac{8}{x}+10"

"f(x)=g(x)=>-2(x+5)+3=2x-5"

"-2x-10+3=2x-5"

"4x=-2"

"x=-0.5"

"g(-0.5)=2(-0.5)-5=-6=f(-0.5)"

Check


"h(-0.5)=\dfrac{8}{-0.5}+10=-6=f(-0.5)=g(-0.5)"

Point "(-0.5, -6)."